LISTSERV mailing list manager LISTSERV 16.5

Help for HEANET-FAULTS Archives


HEANET-FAULTS Archives

HEANET-FAULTS Archives


HEANET-FAULTS@LISTSERV.HEANET.IE


View:

Message:

[

First

|

Previous

|

Next

|

Last

]

By Topic:

[

First

|

Previous

|

Next

|

Last

]

By Author:

[

First

|

Previous

|

Next

|

Last

]

Font:

Proportional Font

LISTSERV Archives

LISTSERV Archives

HEANET-FAULTS Home

HEANET-FAULTS Home

HEANET-FAULTS  April 2005

HEANET-FAULTS April 2005

Subject:

HEA-NOC/20050413-3 [CLOSE] Network wide IOS upgrade

From:

Brian Nisbet <[log in to unmask]>

Reply-To:

Brian Nisbet <[log in to unmask]>

Date:

Tue, 26 Apr 2005 16:50:01 +0100

Content-Type:

text/plain

Parts/Attachments:

Parts/Attachments

text/plain (149 lines)

Ticket Number: HEA-NOC/20050413-3      Ticket Status: CLOSE
Ticket Type:   scheduled               Resolver:      HEA-NOC
Ticket Opened: 20050413 10:32 UTC+1    Problem Start: 20050414 08:00 UTC+1
Ticket Closed: 20050426 16:50 UTC+1    Problem End:   20050425 09:00 UTC+1

Site/line:     Citywest PoP, Cork PoP, DCU-POP PoP, Dublin PoP, Galway PoP,
               Kilcarbery Park PoP, Limerick PoP

Problem Description:
A TCP/IP vulnerability has recently been released.  Exploitation of this
issue could degrade or effectively cease traffic flow between sites on the
network.



For more information on this vulnerability please see:



http://www.niscc.gov.uk/niscc/docs/al-20050412-00308.html?lang=en



The recommended fix for effected systems is to upgrade the Cisco IOS or
Juniper JUNOS.

Problem Effects:
This issue could degrade connectivity or cause some applications that use
TCP across the network to time out.  The Juniper M20s in Cork and New York
are running versions of JUNOS not affected by this exploit.  However
multiple Cisco systems will need to be upgraded.  This will require a
reload of several routers across the network and brief (c. 10 mins)
customer outages will be unavoidable as PoP equipment is upgraded.  It is
envisaged that these reloads will take place over the next week and a
timetable will be circulated.

Problem History:
20050413 10:14 BN Currently identifying correct upgrade path for affected
systems and compiling timeline for upgrades.  This work will start at
08:00 on Thursday 14th April.  Sites that will be affected will be
contacted as soon as possible.

20050413 14:47 BN A timeline has been organised to minimise customer
impact.  The upgrades will take place over the next week, starting
tomorrow at 8am.

Thursday 14th April 08:00 - 10:00
There should be no customer impact, although some connections may be
degraded earlier in the morning and customer connectivity should be
considered at risk.
Outward connectivity to Geant, Janet and Global crossing will lose
redundancy
during this window.  All of these circuits will see BGP peerings go up and
down.
Connection will also be lost to our New York PoP while Deimos.hea.net is
reloaded.  This will affect connectivity to Sinet, CAnet, Abilene, MANLAN
and Startap. 

Hyperion - Brooklawn House PoP Core Router
Deimos - Citywest PoP Core Router
Luna - Kilcarbery Park Core Router
Phobos - Citywest PoP Core Router
Alexander - Old Dub-Gal redundant link - Shut down
Hannibal  - Old Dub-Gal redundant link - Shut down

Monday 18th April 08:00 - 09:00
There should be no customer impact, although some connections in Limerick
and Galway may be degraded earlier in the morning.  This connections
should be considered at risk.

Rhea - UL PoP Core Router
Titan - NUIG PoP Core Router

Tuesday 19th April 08:00 - 09:00
Dublin Customers will have a brief outage during these upgrades.  The
outages are expected to last no more than 10 minutes and will happen prior
to 9am.

Andel - Dublin sub-2Mb serial router
Mantova - Dublin atm/ethernet access router

Wednesday 20th April 08:00 - 10:00
Customers in Cork, Limerick and Galway will have a brief outage during
these upgrades.  The outages are expected to last no more than 10 minutes
and will happen prior to 10am.  HEAnet will endeavour to reload each
router as early as possible during this window.

Leabhair - Cork customer access router
Tempus   - Limerick customer access router
Lucien   - Galway customer access router
Lilith   - INEX v4 router

Thursday 21th April 08:00 - 10:00
Hermes and RIFTTFS customers will have a brief outage during these
upgrades.  DCU's primary circuit will also be down.  The outages are
expected to last no more than 10 minutes and will happen prior to 10am. 
HEAnet will endeavour to reload each router as early as possible during
this window.

Kuiper - DCU PoP access router
Spike - TCD PoP access router
Schiphol - IPv6 Tunnel router
Olympus - Experimental router

20050414 10:10 BN Maintenance Wednesday 14th completed without undue
outage.  Four routers (Phobos, Deimos, Hyperion & Luna) upgraded as
planned.  Old 34Mb Citywest - Galway link shutdown.

20050418 08:42 AH Rhea and Titan (Limerick and Galway core) upgraded to
plan. Traffic observed to reroute via ATM UBRs. Galway reloaded
08:08-08:11 and Limerick  08:31-08:33.

20050419 08:26 OMG Mantova and Andel upgraded according to plan. Mantova
reloaded at 07:56 - 08:01. Andel 08:53 - 08:57.

20050420 08:40 SOB Maintenance for wed 20th completed.  Upgrade for lilith
postponed due to lack of space in the memory for the new IOS.  leabhair,
lucien and tempus all returned to working order and all alarms cleared. 
However the command sh isis neighbor no longer workes in the new IOS.

20050421 09:23 BN All work went as planned this morning.  Olympus
(experimental router) was not reloaded due to other ongoing issues. 
Kuiper, Spike and Schiphol were reloaded sucessfully.  However due to
other issues the three PoP access routers in Cork, Limerick and Galway
will need to be reloaded before 9am on Monday  25th April.  This will
cause an outage to UCC, NCRI, UL, NUIG, CAO and St Angelas.  These outages
will last no more than five minutes.  Affected institutions will be
contacted separately.

20050421 10:44 COC contacted the main technical contacts for CAO, NUIG and
St Angela's College by phone to inform them that the outage would occur.
None of the contacts reported any issues with this planed outage.

20050425 08:44 DC All three PoP access routers reloaded successfully.
Monitoring.

20050425 16:41 BN As all scheduled routers have been reloaded this work is
considered complete.  Situation will be monitored for a further twenty-four
hours before closing ticket.

20050426 16:50  Auto Closed after 1 day.


Time to Fix:
11 days

-- 
This ticket can be monitored at http://www.hea.net/tickets/20050413-3

Top of Message | Previous Page | Permalink

Advanced Options


Options

Log In

Log In

Get Password

Get Password


Search Archives

Search Archives


Subscribe or Unsubscribe

Subscribe or Unsubscribe


Archives

July 2019
June 2019
May 2019
April 2019
March 2019
February 2019
January 2019
December 2018
November 2018
October 2018
September 2018
August 2018
July 2018
June 2018
May 2018
April 2018
March 2018
February 2018
January 2018
December 2017
November 2017
October 2017
September 2017
August 2017
July 2017
June 2017
May 2017
April 2017
March 2017
February 2017
January 2017
December 2016
November 2016
October 2016
September 2016
August 2016
July 2016
June 2016
May 2016
April 2016
February 2016
January 2016
December 2015
November 2015
October 2015
September 2015
August 2015
July 2015
June 2015
May 2015
April 2015
March 2015
February 2015
January 2015
December 2014
November 2014
October 2014
September 2014
August 2014
July 2014
June 2014
May 2014
April 2014
March 2014
February 2014
January 2014
December 2013
November 2013
October 2013
September 2013
August 2013
July 2013
June 2013
May 2013
April 2013
March 2013
February 2013
January 2013
December 2012
November 2012
October 2012
September 2012
August 2012
July 2012
June 2012
May 2012
April 2012
March 2012
February 2012
January 2012
December 2011
November 2011
October 2011
September 2011
August 2011
July 2011
June 2011
May 2011
April 2011
March 2011
February 2011
January 2011
December 2010
November 2010
October 2010
September 2010
August 2010
July 2010
June 2010
May 2010
April 2010
March 2010
February 2010
January 2010
December 2009
November 2009
October 2009
September 2009
August 2009
July 2009
June 2009
May 2009
April 2009
March 2009
February 2009
January 2009
December 2008
November 2008
October 2008
September 2008
August 2008
July 2008
June 2008
May 2008
April 2008
March 2008
February 2008
January 2008
December 2007
November 2007
October 2007
September 2007
August 2007
July 2007
June 2007
May 2007
April 2007
March 2007
February 2007
January 2007
December 2006
November 2006
October 2006
September 2006
August 2006
July 2006
June 2006
May 2006
April 2006
March 2006
February 2006
January 2006
December 2005
November 2005
October 2005
September 2005
August 2005
July 2005
June 2005
May 2005
April 2005
March 2005
February 2005
January 2005
December 2004
November 2004
October 2004
September 2004
August 2004
July 2004
June 2004
May 2004
April 2004
March 2004
February 2004
January 2004
December 2003
November 2003
October 2003
September 2003
August 2003
July 2003
June 2003
May 2003
April 2003
March 2003
February 2003
January 2003
December 2002
November 2002
October 2002
September 2002
August 2002
July 2002
June 2002
May 2002
April 2002
March 2002
February 2002
January 2002
December 2001
November 2001
October 2001
September 2001
August 2001
July 2001
June 2001
May 2001
April 2001
March 2001
February 2001
January 2001
December 2000
November 2000
October 2000
September 2000
August 2000
July 2000
June 2000
May 2000
April 2000
March 2000
February 2000
January 2000
December 1999
November 1999
October 1999
September 1999
August 1999
July 1999
June 1999
May 1999
April 1999
March 1999
February 1999
January 1999
December 1998
November 1998
October 1998
September 1998
August 1998
July 1998
June 1998
May 1998
April 1998
March 1998
February 1998
January 1998
December 1997
November 1997
October 1997
September 1997
August 1997
July 1997
June 1997
May 1997
April 1997
March 1997
February 1997
January 1997
December 1996
November 1996
October 1996
September 1996
August 1996
July 1996
June 1996
May 1996
April 1996
March 1996
February 1996
January 1996
December 1995
November 1995
October 1995
September 1995
August 1995
July 1995
June 1995
May 1995
April 1995
March 1995
February 1995
January 1995
December 1994
November 1994
October 1994
September 1994
August 1994
July 1994
June 1994
May 1994
April 1994
March 1994
February 1994
January 1994
December 1993
November 1993
October 1993
September 1993
August 1993
July 1993
June 1993

ATOM RSS1 RSS2



LISTSERV.HEANET.IE

Secured by F-Secure Anti-Virus CataList Email List Search Powered by the LISTSERV Email List Manager